# Diversity and stylization of the contemporary user-generated visual arts in the complexity-entropy plane

### v2

@@ 3 Discussion
Beyond computationally observing macroscopic evolution of visual art, opportunities for analytically mapping its latent dynamics in a creativity framework remain. Notably, Sigaki et al. [13] has performed a seminal work that demonstrates the feasibility of using the C-H plane to intuitively map the diachronic evolution of visual art styles. The study has recognized “a natural law in the same way as physical growth (Wölfflin)” when affiliating historical alterations in style, particularly from linear to painterly. This remark is similar to David Bohm’s belief about artistic creativity being related to “a harmony parallel to that of nature (Cézanne) [31].”
Apart from the findings of Sigaki et al.’s study providing useful insights into the evolution of visual art styles and its potential relationship to natural laws and creativity, we take an alternative strategy to explore the evolutionary dynamics of creative visual arts in the C-H plane. We look at the evolutionary process underpinning a stylization through the lens of selection, in which a variety of competing agents interact. This state arises by the emergence of novel and different mutations during the reproduction process [32]. A stylization of visual arts entails development of popularization of diverse artistic forms influenced by novel stylistic canons. Therefore, considering diversity as a prerequisite for the emergence of creativity and subsequent stylization, we investigate the C-H plane to empirically identify the optimal C-H condition, by which stylistic diversity is most pronounced while visual artworks of a specific timeframe and groups continue migrating towards homogenic styles.
We find that as the average entropy H rises, so does the intragroup diversity. Our analysis indicate that a novel style emerges in C-H regions where random and diverse agents coexist with strong individual innovativeness (Fig. 1). This finding echoes Eric Hobsbawm, who observed that during the period of highly unpredictable social and technological changes (i.e., the Avant Garde prior to 1914), modern art (low C and high H; Fig. 1a) was “not to claim that it displaced the classic and the fashionable, but that it supplemented both [3].”
@@
Possible limitations of our study include the possibility of selection bias in our data and methodologies for measuring image similarity. Although our dataset contains works by artists across the world, it is dominated by Western artists and is primarily composed of two-dimensional visual artworks. We are also aware that the curated user-generated visual artworks on DeviantArt and Bēhance are likely to represent merely a small portion of a colossal evolution occurring in contemporary visual art disciplines. Regarding our image similarity measures, despite the fact that the ResNet-18 architecture and SIFT capture multi-level image features, there are numerous other possible combinations of methods that capture various spatial image characteristics. Comparing diversity and creativity across different spaces of representation could be an intriguing future research topic.
@@
Future research could investigate the properties of images generated by artificial intelligence (AI) models. In recent years, state-of-the-art AI models (e.g., DALL-E2 [33], stable diffusion models [34], etc.) have made enormous strides in text-to-image art generation. With growing interest in AI-generated images, the relationship between AI and human creativity has come to the forefront of research. A recent study of images generated by a stable diffusion model revealed that AI-generated images tend to have a narrower distribution of entropy and complexity when compared to human drawings; this may lead to less diversity and creativity in their visual artistic expression [35]. In addition, another study addressed the difficulties of evaluating interactions between human and computational creativity in online creative ecosystems [36]. Even though the development of AI art is rapidly advancing and attracting a great deal of attention, we cannot overlook the need to investigate and comprehend its characteristics in a more systematic manner. One could pose queries about the potentials of AI art, with a desire to investigate its representations and inspirations. As our methods are readily applicable to any corpus of two-dimensional digital images and their resulting knowledge maps, we believe the framework of this study can facilitate a better understanding of diversity and stylization in the emerging field of AI art.
@@ 4 Data (materials)
We use the 149,780 images with the C-H values (Table SI 1) during the IE obtaining process. However, the computing process of extraction and pairwise comparison of SIFT features between images consumes considerably more time compared to that of the image embeddings through convolutional neural networks (CNN). Therefore, we take two separate subsampling procedures for EDA on spatiotemporal relationships between local information of the C-H space and image diversity. 1) We randomly subsample images for both IE and SIFT similarity measures for the ARMA model (Table SI 3). 2) Considering each C-H bin sample as a population, we additionally use sample size formula (confidence level: 95%, margin of error: 5%) to determine and randomly subsample the minimum number of necessary samples to meet the desired statistical constraints, and to draw proper inferences from respective C-H bins in Fig. 3.
@@ 5.1 Complexity-entropy (C-H) measures of visual artwork images
Sigaki et al. [13] observed that the change in complexity (C) and entropy (H) of paintings over time could reflect the evolution of art historical styles. Here, we explain how a two-dimensional image is represented in the C-H space. The normalized permutation entropy H and statistical complexity C of an image are calculated based on its matrix representation. First, we investigate submatrices with ordinal patterns for each image with dx by dy embedding dimensions. The normalized Shannon entropy H(P) is then computed with the estimated probability distribution P, which is the probability of the order of neighboring pixel values in an image based on all ordinal patterns within each submatrix for a given embedding parameter (in our case, dx = dy = 2).
$H(P)=\frac{1}{\ln(n)}\sum_{i=0}^{n}p_{i}\ln\left(1/p_{i}\right)$ (1)
The value H signifies the degree of “disorder” in the configuration of the pixels of an image represented by its matrix representation. Moreover, the statistical complexity C(P) is calculated for investigating the degree of structural complexity present in each submatrix,
@@
[…]
We note that in theory, any input image can be plotted as a point in the possible C-H domain bounded by a lower (minimum) and an upper (maximum) parabolic complexity curves, depending on the size of ordinal patterns (based on the embedded parameter (dx = dy = 2)) among the image pixels [38]. Specifically, the Jensen-Shannon divergence between a probability distribution P of local order patterns and the uniform distribution U forms a parabolic range as it is plotted against the given H value [13, 39, 40].
In practice, we calculated the C-H values of artworks using the Ordpy module [41], a simple and open-source Python module that implements permutation entropy and several principle methods for analyzing time series and two-dimensional data using complexity parameters [16, 42, 43].
@@ 5.2 Statistical regression of visual artwork image similarities
Among typical statistical models for time series analysis (e.g., the Bayesian network, the hidden Markov model, and etc.), an autoregressive moving-average (ARMA) model has been used to make inferences of relationships between variables by the form of regression. Parameters of the model are p (order of the autoregressive part specifying the number of lags used) and q (order of the moving average part representing the error of the model as a combination of previous error terms) [44].
We use the ARMA model to regress the variance of the similarities among the artworks on the degree of entropy H, complexity C, and their variances over the years of 2010 to 2020 (Table 1). The Modified Dickey-Fuller unit-root test is used to determine the number of lags required to achieve a stationary series [45], resulting in 1 lag for SIFT models, and 3 lags for IE models. Additionally, the mean and variance of the entropy H and complexity C are calculated using all the images from respective years.
@@ 5.3 Multi-level image features for visual artwork similarity measures
Recent analyses of visual art styles use style vectors based on deep CNN to extract stylistic - i.e., high-level features and information from paintings [46]. On the other hand, the SIFT feature descriptor as a well-established CV technique has been demonstrated that the descriptor is effective for a variety of objectives related to image matching and object recognition. CNN filters by themselves perform similarly to SIFT in detecting low-level and straightforward - i.e., hand-crafted local invariant features, while outperforming SIFT in detecting high-level and complex features [47]. The combined usage of CNN and SIFT features has proven to result in discriminative multi-level image features using the best of both [48].
We initially build on ResNet-18 architecture pretrained with the ImageNet database to construct multi-level image features in our image set. ResNet pretrained with ImageNet has also been shown to provide sufficient feature representations for paintings [19], making it appropriate for extracting latent visual features - i.e. image embeddings (IE) - from heterogeneous visual artworks.
Inspired by Liu et al.’s approach of extracting multi-level features of artworks to represent art style of images [21], we obtain the embeddings of our images in the following manner (see Fig. 2b for the detailed technical pipeline). 1) As combining both low- and high-level features are effective for encoding art styles, the max (front) and average (back) pooling layers of ResNet-18 are chosen as the convolutional layers to extract hidden feature maps from each input image (224 × 224 × 3). 2) We then reduce dimensionality of the hidden low-level feature map (56 × 56 × 64) extracted from the maximum pooling layer to a one-dimensional feature map (56 × 56 × 1). 3) We also extract a hidden high-level feature map (1 × 512) from the average pooling layer just prior to the fully connected layer, where no feature map exists. 4) We then run a Principal Component Analysis (PCA) on both low-level and high-level feature maps to derive 100-dimensional embedding from each, thereby creating a 200-dimensional embedding vector of each user-generated artwork. 5) Lastly, we measure pairwise cosine similarities between embeddings of the contemporary user-generated visual artworks.
@@
The data underlying the analyses and findings of this study will be available from S.K. (ryankim990@gmail.com) on a reasonable request. Researchers interested in accessing data can contact the corresponding author (wnjlee@kaist.ac.kr). Correspondence and requests for data should be addressed to S.K. and W.L.

### v3

@@ 3 Discussion
Beyond computationally observing the macroscopic evolution of visual art, opportunities for analytically mapping its latent dynamics in a creativity framework remain. Notably, Sigaki et al. [13] have performed a seminal work that demonstrates the feasibility of using the C-H plane to intuitively map the diachronic evolution of visual art styles. The study has recognized “a natural law in the same way as physical growth (Wölfflin)” when affiliating historical alterations in style, particularly from linear to painterly. This remark is similar to David Bohm’s belief about artistic creativity being related to “a harmony parallel to that of nature (Cézanne) [31].”
This study quantitatively tracks user-generated artworks of the contemporary visual art disciplines within the C-H space and interprets their stylization from the perspective of diversity within groups. Apart from the findings of Sigaki et al.’s study providing useful insights into the evolution of visual art styles and their potential relationship to natural laws and creativity, we take an alternative strategy to explore the evolutionary dynamics of creative visual arts in the C-H plane. We look at the evolutionary process underpinning a stylization through the lens of selection, in which a variety of competing agents interact. This state arises from the emergence of novel and different mutations during the reproduction process [32]. A stylization of visual arts entails the development of popularization of diverse artistic forms influenced by novel stylistic canons. Therefore, considering diversity as a prerequisite for the emergence of creativity and subsequent stylization, we investigate the C-H plane to empirically identify the optimal C-H condition, by which stylistic diversity is most pronounced while visual artworks of a specific timeframe and groups continue migrating towards homogenic styles.
We find that as the average entropy H rises, so does the intragroup diversity. Our analyses indicate that a novel style emerges in C-H regions where random and diverse agents coexist with strong individual innovativeness (Fig. 1). This finding echoes Eric Hobsbawm, who observed that during the period of highly unpredictable social and technological changes (i.e., the Avant Garde prior to 1914), modern art (low C and high H; Fig. 1a) was “not to claim that it displaced the classic and the fashionable, but that it supplemented both [3].”
@@
Building on the study by Sigaki et al. [13], our research provides a novel perspective on contemporary visual art disciplines by positioning user-generated artworks within the complexity-entropy (C-H) space. This framework quantifies stylization while revealing diversity and coherence within artistic groups. Our work extends the analysis of C-H trajectories of visual art stylization from the early 20th century to the 1970s, confirming an intensified tendency of contemporary visual artworks to shift toward the upper-left region within the given C-H framework. By situating our findings within art history, parallels can be drawn with earlier work that used neural networks to trace artistic influences across historical movements [33], as both approaches highlight how stylistic diversity reflects creative innovation and adherence to group norms. Similarly, our entropy-based analysis aligns with a study that applied spatial entropy to landscape analysis [34], underscoring entropy’s utility in analyzing structural and compositional diversity across different contexts.
Our research also complements the findings of Valensise et al. [35], who used the C-H plane to analyze meme evolution, highlighting how community-driven curation fosters a balance between novelty and conformity. This dynamic is evident in curated subsets like DeviantArt’s “Daily Deviation” and Bēhance’s “Best of Bēhance,” which mirror the evolutionary patterns of digital culture. Additionally, our work aligns with Deng et al. [36], who examined representativity in art paintings. Their study underscores the challenges of quantifying art’s diversity while preserving its nuanced cultural significance. By extending these efforts to contemporary digital art, we bridge historical and modern perspectives, demonstrating how user-generated art mirrors broader trends in collective artistic evolution.
Meanwhile, our study has several limitations, including potential selection bias in our data and the methodologies used for measuring image similarity. While our dataset includes works by artists from around the world, it is predominantly composed of two-dimensional visual artworks created by Western artists. Furthermore, the curated user-generated artworks on DeviantArt and Bēhance during the specified time period likely represent only a fraction of the vast and dynamic evolution occurring within contemporary visual art disciplines.
As platforms showcasing global creativity, their curated collections are expected to result in cultural and social diversity. However, many independently operated platforms within specific cultural contexts remain unexplored, presenting valuable opportunities for further research. Future studies could address these limitations by incorporating a broader range of data sources, including other social media and regional art-sharing platforms, to enhance the representation of global artistic and cultural diversity, and to validate and expand upon our findings.
Regarding our image feature extraction methods, despite the fact that the ResNet-18 architecture and SIFT capture multi-level image features, there are numerous other possible combinations of methods that capture various spatial image characteristics. Understanding the limitations of our image similarity measurement methods and how they impact our analyses is crucial. The SIFT algorithm is sensitive to image variations (e.g., significant affine transformations, illumination changes, and image noise) possibly affecting the detection and description of keypoints. Also, possible loss of subtle yet important features or fine-grained details cannot be ignored due to reduced dimensionality in ResNet latent PCA embeddings. Therefore, exploring alternative or complementary visual features in future experiments could provide more robust and generalizable insights.
The rapid advancements in computer vision have introduced a variety of image processing models, each offering unique capabilities for analyzing visual data. Different choices of feature extraction models may produce moderately varying similarity outcomes, thereby shaping diversity assessments and influencing our understanding of the evolution of art styles. Notably, transformer-based models, such as Vision Transformers (ViTs) [37], have demonstrated potential for capturing global context across entire images, providing richer and more nuanced feature embeddings. Specifically, CLIP (Contrastive Language-Image Pretaining) [38] and DINO (Distillation of Non-contrastive Image Representations) [39] provide embeddings aligned with semantic meaning, which could improve high-level similarity detection. While our research focused on assessing the “diversity of visual forms” by focusing more on visual information itself, future studies could explore the diversity and evolution of contemporary visual arts from a new perspective. This could involve utilizing state-of-the-art models to measure artwork similarity based on semantic and contextual information, offering a deeper understanding of artistic trends and styles.
@@
Future research could also investigate the properties of images generated by artificial intelligence (AI) models. In recent years, state-of-the-art AI models (e.g., DALL-E2 [40], stable diffusion models [41], etc.) have made enormous strides in text-to-image art generation. With growing interest in AI-generated images, the relationship between AI and human creativity has come to the forefront of research. A recent study of images generated by a stable diffusion model revealed that AI-generated images tend to have a narrower distribution of entropy and complexity when compared to human drawings; this may lead to less diversity and creativity in their visual artistic expression [42]. In addition, another study addressed the difficulties of evaluating interactions between human and computational creativity in online creative ecosystems [43]. Even though the development of AI art is rapidly advancing and attracting a great deal of attention, we cannot overlook the need to investigate and comprehend its characteristics in a more systematic manner. One could pose queries about the potentials of AI art, with a desire to investigate its representations and inspirations. As our methods are readily applicable to any corpus of two-dimensional digital images and their resulting knowledge maps, we believe the framework of this study can facilitate a better understanding of diversity and stylization in the emerging field of AI art.
@@ 4 Data (materials)
Meanwhile, the vast size of the entire collection of user-generated artworks on these platforms makes comprehensive analysis impractical. Therefore, we processed a more manageable and meaningful dataset of high-quality artworks broadly recognized as influential within their respective artistic communities. To ensure that the images analyzed represented broad trends in contemporary user-generated visual art, we specifically focused on subsets of artworks curated by the platforms themselves: “Daily Deviation” on DeviantArt and “Best of Bēhance” on Bēhance. These subsets are quasi-canonical, as they are promoted daily by the platforms based on quality, innovation, and relevance, reflecting current trends in user-generated art.
We use the 149,780 images with the C-H values (Table SI 1) during the IE obtaining process. However, the computing process of extraction and pairwise comparison of SIFT features between images consumes considerably more time compared to that of the image embeddings through convolutional neural networks (CNN). Therefore, we take two separate subsampling procedures for EDA on spatiotemporal relationships between local information of the C-H space and image diversity. 1) We randomly subsample images for both IE and SIFT similarity measures for the ARMA model (Table SI 3). 2) Considering each C-H bin sample as a population, we additionally use the sample size formula (confidence level: 95%, margin of error: 5%) to determine and randomly subsample the minimum number of necessary samples to meet the desired statistical constraints, and to draw proper inferences from respective C-H bins in Fig. 3.
@@ 5.1 Complexity-entropy (C-H) measures of visual artwork images
Sigaki et al. [13] observed that changes in complexity (C) and entropy (H) of paintings over time could reflect the evolution of art historical styles. Here, we explain how a two-dimensional image is represented in the C-H space. We calculate the normalized permutation entropy H and statistical complexity C of an image based on its matrix representation. The original image files are in JPEG and PNG formats, represented in a 24-bit RGB color space (8 bits each for the red, green, and blue channels). We obtain a simpler matrix representation of each image by averaging the three color channels of every pixel.
Next, we examine submatrices with ordinal patterns for each image using embedding dimensions $d_{x}\times d_{y}$ (in our case, $d_{x}=d_{y}=2$), leading to ($(d_{x}d_{y})!=24$) possible ordinal patterns. By sliding partitions of size $d_{x}\times d_{y}$ pixels across the entire matrix, we obtain a distribution of ordinal patterns $P$ for the image. Here, $P=\{p_{i};i=1,\dots,n\}$ can be viewed as a 24-dimensional vector whose elements sum to 1.
From the probability distribution $P$, we calculate the normalized Shannon entropy $H(P)$:
$H(P)=\frac{1}{\ln(n)}\sum_{i=1}^{n}p_{i}\ln\left(1/p_{i}\right)$ (1)
, where $n=(d_{x}d_{y})!$.
The entropy H approaches 1 if the pixel order appears random, and $H\approx 0$ if the pixel order is highly regular and always appears in the same configuration. The value of H signifies the degree of ‘disorder’ in the configuration of the pixels in an image, as represented by its matrix representation.
Furthermore, $C(P)$ is calculated to investigate the degree of structural complexity present in the submatrices. $C(P)$ is defined as:
@@
, where $U=\{u_{i}=1/n\;;\;i=1,\dots,n\}$ represents the uniform distribution, $D(P,U)$ is the Jensen-Shannon divergence between $P$ and $U$, and $D^{*}$ is the maximum Jensen-Shannon divergence.
The Jensen-Shannon divergence $D(P,U)$ is given by:
$D(P,U)=S\left(\frac{P+U}{2}\right)-\frac{S(P)}{2}-\frac{S(U)}{2}$ (3)
, where $(P+U)/2=\left\{(p_{i}+1/n){2},\;i=1,\dots,n\right\}$. The maximum divergence $D^{*}$ can be calculated as:
$D^{*}=\max D(P,U)=\frac{1}{2}\left[\frac{n+1}{n}\ln(n+1)+\ln(n)-2\ln(2n)\right]$ (4)
This is obtained when $P=\{p_{i}=\delta_{1,i}\;;\;i=1,\dots,n\}$, where $\delta_{1,i}$ is the Kronecker delta.
$C(P)$ increases as the distribution $P$ of the local order patterns deviates from the uniform distribution $U$ or as the entropy of $P$ increases. Specifically, $C(P)$ reaches its maximum in a state that is neither completely uniform nor entirely homogeneous. Stylistically, $C(P)$ reflects how much the objects within an image are spatially circumscribed or interrelated, while $H(P)$ reflects how distinctly the objects are outlined or how fluidly they are intertwined [15].
In practice, we calculate the C-H values of artworks using the Ordpy module [45], a simple and open-source Python module that implements permutation entropy and several principle methods for analyzing time series and two-dimensional data using complexity parameters [16, 46, 47]. To obtain the upper- and lower-boundary curves in the C-H plane for our 2D images, we employed two functions maximum_complexity_entropy and minimum_complexity_entropy in the Ordpy module [45]. These functions systematically generate multiple probability distributions over the possible ordinal patterns and compute both $H(P)$ and $C(P)$. They do so by assigning probability mass to one or a small number of ordinal patterns while distributing the remaining probability evenly among the other patterns, scanning through this space of distributions for H and C values. Collecting and sorting the H and C points yields approximate upper- and lower-boundary curves that envelop the values found in experimental or simulated data, thus delineating the limits in the C-H plane.
Lastly, for the choice of embedding dimensions, we set $d_{x}=d_{y}=2$. While these values are tuning parameters that can be adjusted, it is known from previous studies that $(d_{x}d_{y})!$ $\ll$ (width × height) must hold to ensure reliable results [16]. Given that the average image width and height of 149,780 images used in this study are approximately 1,000 pixels (average width: 1011.61px, average height: 953.91px; refer to Table SI 1), we chose $d_{x}=d_{y}=2$ as the embedding dimensions.
@@ 5.2 Statistical regression of visual artwork image similarities
Among typical statistical models for time series analysis (e.g., the Bayesian network, the hidden Markov model, etc.), we used the ARMA model to regress the dyadic similarities among the artworks on the degree of complexity C, entropy H, and their variances over the respective years from 2010 to 2020 (Table 1):
$y_{t}=X_{t}\beta+\sum_{i=1}^{p}\phi_{i}y_{t-i}+\sum_{i=1}^{q}\theta_{i}\varepsilon_{t-i}+\varepsilon_{t}$ (5)
, where $y_{t}$ is the average value of the dyadic similarities in DeviantArt and Bēhance at year $t$, and $X_{t}$ is a vector of the C-H covariates at $t$. $p$ and $q$ are the number of lags for autocorrelation and moving-average, respectively. In order to obtain a stationary series of our data, we used the modified Dickey-Fuller unit-root test [48], which determined the parameter $p$. This results in $p=3$ for IE models, and $p=1$ for SIFT models. $q$ was determined to be 1 for both IE and SIFT models using Bartlett’s approximation [49].
@@ 5.3 Multi-level image features for visual artwork similarity measures
Recent analyses of visual art styles use style vectors based on deep CNN to extract stylistic (i.e., high-level) features and information from paintings [50]. On the other hand, the SIFT feature descriptor as a well-established CV technique has been demonstrated that the descriptor is effective for a variety of objectives related to image matching and object recognition. CNN filters by themselves perform similarly to SIFT in detecting low-level and straightforward (i.e., handcrafted) local invariant features, while outperforming SIFT in detecting high-level and complex features [51]. The combined usage of CNN and SIFT features has proven to result in discriminative multi-level image features using the best of both [52].
We initially build on ResNet-18 architecture pretrained with the ImageNet database to construct multi-level image features in our image set. ResNet pretrained with ImageNet has also been shown to provide sufficient feature representations for paintings [19], making it appropriate for extracting latent visual features - i.e. image embeddings (IE)—from heterogeneous visual artworks.
Inspired by Liu et al.’s approach [21] of extracting multi-level features of artworks to represent art styles of images, we obtain the embeddings of our images in the following manner (see Fig. 2b for the detailed technical pipeline). 1) As combining both low- and high-level features is effective for encoding art styles, the max (front) and average (back) pooling layers of ResNet-18 are chosen as the convolutional layers to extract hidden feature maps from each input image (224 × 224 × 3). 2) We then reduce the dimensionality of the hidden low-level feature map (56 × 56 × 64) extracted from the maximum pooling layer to a one-dimensional feature map (56 × 56 × 1). 3) We also extract a hidden high-level feature map (1 × 512) from the average pooling layer just prior to the fully connected layer, where no feature map exists. 4) We then run a Principal Component Analysis (PCA) on both low-level and high-level feature maps to derive 100-dimensional embedding from each, thereby creating a 200-dimensional embedding vector of each user-generated artwork. 5) Lastly, we measure pairwise cosine similarities between embeddings of the contemporary user-generated visual artworks.
@@
The data underlying the analyses and findings of this study will be available from S.K. (ryankim1101@kaist.ac.kr) on a reasonable request. Researchers interested in accessing data can contact the corresponding author (wnjlee@kaist.ac.kr). Correspondence and requests for data should be addressed to S.K. and W.L.
@@
Acknowledgements
We would like to thank Lev Manovich, Kangsan Lee, and Frédéric Godart for their valuable feedback and constructive suggestions on earlier versions of this manuscript. We sincerely thank the anonymous reviewers for their helpful comments to improve this manuscript. S.K., B.L., and W.L. acknowledge the support from the BK21 Plus Postgraduate Organization for Content Science (N20210014).
Supplementary Information
Figures and tables
Platform Data type Initial Measurement Uncomputable Zero C-H Processed Filtered Image
samples samples samples samples samples dimensions
DeviantArt 70,804 210 4 70,590 68,735 Width (avg.): 1013.15px,
JPEG Complexity-entropy Height (avg.): 1023.82px
and (C-H) values
PNG per image Width (avg.): 1010.31px,
Bēhance 348,040 9,939 2,685 335,416 81,045 Height (avg.): 894.61px
Total. 418,844 10,149 2,689 406,006 ($\approx$ 97%) 149,780 ($\approx$ 36%) $\approx$ 1,000px
Table SI 1: Dataset description for complexity-entropy (C-H) measure. We directly downloaded user-generated visual artwork images via image URLs extracted from the web pages of the corresponding projects that have been curated and promoted on a daily basis between early 2010 and the end of 2020 in DeviantArt (Daily Deviations) and Bēhance (Best of Bēhance). Following the extraction of JPEG and PNG images from our initial image set for image quantification, our initial sample consists of a total of 418,844 artworks, including 70,804 images from DeviantArt and 348,040 images from Bēhance. Subsequently, we generate annual time series data for empirical studies in this paper by combining all the image data values into yearly sample datasets based on corresponding platforms and the years of the original projects’ showcase dates. As for the C-H measures, we exclude 10,149 images that could not be computed by the cv2 module as ‘abnormal’ sets, as well as 2,689 images with both C-H values of 0. The rest of the 406,006 images with respective C and H values are then filtered to 149,780 images of sub-visual art fields of DeviantArt and Bēhance. Given that the average image width and height of 149,780 images used in this study are approximately 1,000 pixels (average width: 1011.61px, average height: 953.91px), we chose $d_{x}=d_{y}=2$ as the embedding dimensions.
Figure SI 1: Cross validating predictive accuracies of statistical learning algorithms. For a robustness check, we cross-validate the predictive accuracy of C-H values from our dataset using four machine learning algorithms (nearest neighbors, random forest, support vector machine, and neural network). We consider the classification of C-H values of diverse artworks based on their original platforms and the year of curation. As a result, it is confirmed that the accuracy scores of the overall classifiers are considerably higher than those of the dummy classifiers: stratified $\approx$ 10% (DeviantArt), $\approx$ 9.3 - 9.4% (Bēhance), and uniform $\approx$ 9.1 - 9.2% (DeviantArt), 9.3 - 9.4% (Bēhance)). Each platform’s statistical algorithms have comparable mean accuracy (DeviantArt: $\sim$ 15% and Bē$\sim$ 13.7%). In line with cross-validation results regarding C-H measures from Sigaki et al., our results demonstrate that the C-H measures of contemporary visual artworks accurately predict both the years and platforms to which they belong. This implies a descriptive evaluation of the degree to which individual images correspond to each year label in the C-H plane.
Creative fields - Bēhance N of samples (prop.)
3D Art 263 (0.32%)
Architecture 5,562 (6.86%)
Crafts 1,735 (2.14%)
Fashion 4,728 (5.83%)
Fine Arts 6,611 (8.16%)
Game Design 1,355 (1.67%)
Graphic Design 27,591 (34.04%)
Illustration 59,048 (72.86%)
Photography 4,434 (5.47%)
UI/UX / Information Architecture 1,708 (2.11%)
Animation 6,043 (7.46%)
Cartooning 3,091 (3.81%)
Character Design 19,459 (24.01%)
Table SI 2: Number of samples associated with clustered sub-topics of Bēhance—normalized proportion of samples in each creative field over the total number of samples in Bēhance with each sample associated with multiple creative fields. The difference in the overall positions of the two platforms in the C-H plane is due to the different proportions of genres that make up each platform. The majority of the Bēhance image set consists of images from Graphic Design and Illustration, while the DeviantArt image set is mostly composed of paintings featuring pictorial objects. Table SI 1 summarizes the distribution of creative fields in the Bēhance dataset. In the case of DeviantArt, due to the absence of image classification from the individual project pages in the platform, the categorization of images is manually conducted by examining a sample size of 1,000 images. Consequently, 72.1% of the images are drawings or paintings containing pictorial objects, and the proportion of photos and design works is notably small.
Figure SI 2: Five creative fields of Bēhance with the greatest sample proportion and their annual mean C-H values (2010-2020). (a) Yearly trends of sample proportions from the creative fields (Illustration, Graphic Design, Character Design, Animation, and Fine Arts) in Bēhance (2010-2020). Comparing samples from 2010 and 2020, the annual proportion of the Illustration field in 2010 (64.27%, n = 3,491) increase by 4,680 images, accounting for 76.92% (n = 8,171) of samples from 2020. The Graphic Design field’s annual proportion in 2010 (32.36%, n = 1,758) increased by 3,368 images, making it the second highest proportion of the year 2020 samples (48.25%, n = 5,126). The other two categories are Character Design and Animation, which increase by 1,451 images and 939 images, respectively, making them the third (19.95%, n = 2,119) and fourth (9.5%, n = 1,057) proportional fields by 2020. In 2010, the annual percentage and sample size for the Fine Arts field (12.79%, n = 984) decreased to 6.06% (n = 644). (b) Yearly average C-H values of visual artworks from the creative fields (Illustration, Graphic Design, Character Design, Animation, and Fine Arts) in Bēhance (2010-2020). Concerning the five fields’ annual C-H movements, the annual average H values of their samples decrease over a decade, indicating that their styles become more linear. Both the annual average C and H values for Animation and Character Design decrease over time, indicating that they become less complex and more linear. In addition, the Illustration, Graphic Design, and Fine Arts sectors exhibit an increase in their annual average C values over time, indicating that their styles incorporate incremental continuums of visually interconnected elements.
Figure SI 3: Pairwise correlations between image similarity measures. (a, b, c) Pairwise correlations between IE similarity, SIFT similarity, and Euclidean distance (C-H) of image representations in 5,000 pairs of DeviantArt artworks. (d, e, f) Pairwise correlations between IE similarity, SIFT similarity, and Euclidean distance (C-H) of image representations in 5,000 pairs of Bēhance artworks. We verify correlations between image similarity measures used in this paper. Marginal plots with regression lines of two different similarity measures of 5,000 unique pairs of visual art images across the entire C-H plane (C: 0 - 1, H: 0 - 1) altogether show weak correlations (refer to correlation coefficients in each plot) in DeviantArt and Bēhance.
Measurement N of random subsamples Processed pairs
Pairwise cosine similarities of
IE of images
11,000 images per platform
(1,000 images per year)
5,494,500 pairs per platform
(499,500 pairs per year)
Pairwise Jaccard similarities of
SIFT features of images
1,100 images per platform
(100 images per year)
19,800 pairs per platform
(4,950 pairs per year)
Table SI 3: Dataset description of image similarity measures used for the ARMA model. To calculate the average image similarity from the online platforms for each given year, we randomly subsample 1) 1,000 images from each year image set and calculate the mean of the 499,500 pairwise IE similarities, and 2) 100 images from each year image set and calculate the mean of the 4,950 pairwise SIFT similarities.
Figure SI 4: Visualization of spatial relationships between the C-H space and the observed image diversity in WikiArt data as measured by the image dissimilarity. (a, b, c) Degrees of sample density, IE similarity (sample mean), and SIFT similarity (subsample mean) of the WikiArt images per bin in the C-H region. Each C-H bin (C $\approx$ 0.01, H $\approx$ 0.02) within a partitioned C-H region (C: 0 - 0.31, H: 0.5 - 1) per platform occupies more than 10 images.